OpenStudy (anonymous):

I need to convert these to a mixed number in simplest form. I can only get so far, and then I think I am doing it wrong. This is what I have done so far... thanks for your help!! -8 3/4 + -11 1/3 = ?. 12 is the LCD, so I get-29/4 = -34/3. Extended out to 261/12 = 136/12. My final answer is 397/24 which I know is wrong... :(

OpenStudy (judygreeneyes):

-8 3/4 + (-11 1/3) = ? You started out correctly by finding the LCD=12, and I think you converted the fractions into twelfths correctly, but what you should end up with after doing that is this for the two fractions (leaving both of them in twelfths, don't reduce them yet!): -8 3/4 = -(96/12 + 9/12) = -105/12 -11 1/3 = -(132/12 + 4/12) = -136/12 Now you have -105/12 - 136/12 = - 241/12 I think this is as simplified as the fraction gets, or you can write it as - 20 1/12

OpenStudy (anonymous):

thank u Judy!! :)

OpenStudy (judygreeneyes):

Very welcome!